OpenStudy (anonymous):

A phonograph record has an initial angular speed of 34 rev/min. The record slows to 12 rev/min in 1.1 s. What is the record’s average angular accel- eration during this time interval? Answer in units of rad/s2

OpenStudy (jdoe0001):

\(\bf \textit{the phonograph went from }\cfrac{34rev}{min}\textit{ to }\cfrac{12rev}{min}\textit{ what's the difference?}\\ \quad \\ \cfrac{22rev}{min}\quad \textit{let us convert to radians over seconds first}\\ \quad \\ \cfrac{22rev}{min}\cdot \cfrac{min}{60sec}\cdot \cfrac{2\pi}{rev}\implies \cfrac{44\pi}{60sec}\) the average acceleration will be rate of change, over how long it took so, it only took 1.1 second for that change. so \(\bf \cfrac{\frac{44\pi}{60sec}}{1.1sec}\)